### v1

# Decoding Decoded: Understanding Hyperparameter Effects in Open-Ended Text Generation

Abstract
Decoding strategies for large language models (LLMs) are a critical but often underexplored aspect of text generation tasks. Since LLMs produce probability distributions over the entire vocabulary, various decoding methods have been developed to transform these probabilities into coherent and fluent text, each with its own set of hyperparameters. In this study, we present a large-scale, comprehensive analysis of how hyperparameter selection affects text quality in open-ended text generation across multiple LLMs, datasets, and evaluation metrics. Through an extensive sensitivity analysis, we provide practical guidelines for hyperparameter tuning and demonstrate the substantial influence of these choices on text quality. Using three established datasets, spanning factual domains (e.g., news) and creative domains (e.g., fiction), we show that hyperparameter tuning significantly impacts generation quality, though its effects vary across models and tasks. We offer in-depth insights into these effects, supported by both human evaluations and a synthesis of widely-used automatic evaluation metrics.
Decoding Decoded:
Understanding Hyperparameter Effects in Open-Ended Text Generation
Esteban Garces Arias1,2, Meimingwei Li1, Christian Heumann 1, Matthias Aßenmacher 1,2 1Department of Statistics, LMU Munich, 2Munich Center for Machine Learning (MCML) Correspondence: Esteban.GarcesArias@stat.uni-muenchen.de
Figure 1: The quality of machine-generated text, as assessed by both automatic metrics and human judgment, is strongly influenced by hyperparameter selection.
1 Introduction
Large language models (LLM) do not readily output natural language text, but instead provide the user only with a high-dimensional probability distribution. The process of transforming these probabilities into text is referred to as decoding and can have nearly as much of an impact on the quality of the generated text as the LLM backbone itself. Several decoding strategies currently employed in conjunction with contemporary state-of-the-art (SOTA) LLMs encompass hyperparameters playing critical roles and hence significantly influencing the quality of the generated text. However, despite their importance, the selection and tuning of these hyperparameters remain an under-explored area, with the users often resorting to defaults and rather benchmarking different models instead of decoding strategies. In particular, this neglects differences in the desired properties of the generated texts, ranging from factual domains, like news generation, to more creative ones like (fiction) literature. Further, the success of decoding strategies might differ between models, a phenomenon that cannot be uncovered by the current practice of only assessing strategies globally, i.e. across models. Recent research emphasizes that different hyperparameter configurations can greatly impact both the coherence and diversity of generated text, underscoring the need for systematic evaluation Zhou et al. (2024).
In this study, we systematically vary hyperparameters of commonly used decoding strategies to assess their effect on text quality across different data sets and importantly, across different SOTA open-source models. By using pre-defined grids for the hyperparameters in deterministic (beam width), stochastic ($k$, $p$, and temperature), and contrastive ($\alpha$ and $k$) decoding strategies, we explore their role in balancing creativity, trustworthiness, and overall text quality. This is especially pertinent in addressing the common problem of degeneration and ensuring effective adaption to diverse text generation tasks Shi et al. (2024). Moreover, modern LLMs, while highly advanced, are still prone to issues like incoherence or degeneration under various hyperparameter settings. Our work addresses these challenges by providing a sensitivity analysis, offering practical guidelines for tuning decoding hyperparameters to optimize performance while mitigating degeneration
Contributions
This study advances the field of decoding strategies for LLMs by performing a large-scale sensitivity analysis of commonly used methods and examining their practical effects on model performance.
  1. 1.
    We conduct an extensive experimental analysis to evaluate the impact of hyperparameter values on various text quality metrics in open-ended text generation.
  2. 2.
    Our comprehensive analysis reveals key factors that influence the quality of LLM-generated texts, as assessed by widely adopted evaluation metrics covering multiple lexical dimensions.
  3. 3.
    Based on these insights, we offer actionable recommendations for practitioners to select appropriate decoding strategies and hyperparameters tailored to specific use cases.
  4. 4.
    In total, we generate 2.2 million text continuations, which are publicly available for future meta-analyses, along with our complete codebase: https://github.com/YecanLee/Decoding-Decoded
2 Decoding Strategies and Hyperparameters
Decoding strategies are generally categorized into two types: deterministic and stochastic. Given the complexity introduced by their hyperparameters, we additionally separate contrastive strategies (cf. Sec. 2.3) from the other deterministic ones (cf. Sec. 2.1). While the latter is solely focused on maximizing the joint probability, the former explicitly compromise this objective by incentivizing more diverse texts.
2.1 Deterministic Strategies
Deterministic strategies follow fixed decision-making processes that do not incorporate randomness. These methods are widely used in tasks requiring high reproducibility and reliability.
Greedy search.
This strategy selects the token with the highest probability at each time step, yielding a sequence that maximizes the local likelihood. However, it often leads to suboptimal results due to the lack of look-ahead, which can cause the model to get trapped in less coherent output.
Beam search.
Freitag and Al-Onaizan (2017) extends the greedy approach by maintaining a beam of the $w$ most probable sequences at each time step. The hyperparameter $w$, known as the beam width, controls the trade-off between diversity and computation. Beam search reduces the risk of suboptimal choices but can still generate repetitive or generic text in open-ended tasks.
Frustratingly Simple Decoding (FSD).
Yang et al. (2024b) propose a method that contrasts outputs from the LLM with an auxiliary anti-LM, constructed using the current prefix. This method seeks to amplify the distinctions between probable and improbable tokens, yielding more robust generation in certain contexts.
2.2 Sampling Strategies
Sampling-based strategies introduce stochasticity to encourage diversity in text generation. These methods allow for more creative and varied outputs but require careful tuning of hyperparameters to balance coherence and randomness.
Sampling with temperature.
Introduced by Ackley et al. (1985), this method samples from the full token distribution while modifying the softmax input. The temperature hyperparameter controls the sharpness of the distribution—higher temperatures flatten it, increasing randomness, while lower temperatures make it more deterministic, favoring higher-probability tokens.
Top-k Sampling.
Proposed by Fan et al. (2018), this strategy limits sampling to the top $k$ tokens with the highest probabilities, forming a subset $V^{(k)}$. By truncating the distribution, it reduces the risk of sampling from low-probability (and potentially incoherent) tokens, while still introducing some diversity.
Nucleus Sampling.
Also known as $p$-sampling, this method was introduced by Holtzman et al. (2019). It samples from the smallest subset $S$ whose cumulative probability exceeds a threshold $p$. Nucleus sampling adapts dynamically to the token distribution, preserving both diversity and relevance by considering only the most probable tokens.
Typical sampling.
Meister et al. (2023) developed this method to produce more diverse and interesting text by aiming for tokens with an expected level of information content. It avoids both overly probable and highly improbable tokens, resulting in a more balanced generation suited for open-ended contexts.
2.3 Contrastive Strategies
Contrastive strategies leverage comparisons between different models or hypotheses to improve text quality. These methods focus on enhancing coherence and informativeness without sacrificing diversity.
Contrastive Search.
Su et al. (2022) introduce a look-ahead mechanism that penalizes tokens disrupting the isotropy of the latent space in the language model, leading to more coherent and semantically consistent text.
Contrastive Decoding.
Li et al. (2023) propose searching for tokens that maximize the difference in probability between a large language model (LLM) and a weaker auxiliary model, helping to refine generation quality by selecting tokens that best distinguish high-quality text.
Adaptive Contrastive Decoding.
(Zhu et al., 2024) and Adaptive Contrastive Search. (Garces-Arias et al., 2024) propose adaptive strategies that adjust to model uncertainty (measured by Shannon entropy) at each time step. These methods aim to strike a balance between coherence and diversity dynamically, removing the need for extensive hyperparameter tuning.
3 Related Work
The choice of decoding strategy plays a crucial role in shaping the performance of text generation models, significantly affecting the trade-offs between output quality and diversity. Wiher et al. (2022) provide a foundational analysis of decoding methods such as beam search, top-k sampling, and nucleus sampling. Their findings indicate that while beam search is effective for structured tasks like machine translation, it often leads to repetitive or less coherent text in creative tasks like story generation. This highlights the need to adapt decoding strategies to specific task requirements.
Subsequent research, including Amini et al. (2023), builds on these insights by offering a comprehensive review of the principles guiding text generation. The study emphasizes how various decoding mechanisms influence the final output, underscoring the necessity of selecting appropriate strategies based on the task at hand. It also explores the balance between fluency and diversity, particularly in open-ended text generation tasks.
More recently, Shi et al. (2024), investigates how decoding strategies scale with large language models (LLMs). This work explores the interaction between decoding strategies, hyperparameters such as top-k and temperature, and model size. The authors illustrate the increasing complexity of optimizing these strategies as models grow in scale, particularly in open-ended generation scenarios.
Additionally, Zhou et al. (2024) offers an in-depth exploration of sampling-based methods. It provides guidelines on managing the balance between diversity and the risk of incoherence, demonstrating how hyperparameter tuning can shape the quality and diversity of generated outputs. This work serves as a practical resource for selecting decoding strategies based on task-specific needs.
However, most existing studies fail to provide an extensive and detailed analysis of how various quality metrics, such as coherence, diversity, and MAUVE, are affected by different decoding strategies. To address this gap, we introduce a new metric, QText, which combines these metrics using the harmonic mean. We compare the best and worst performing strategies based on QText with those identified by human evaluators, highlighting areas of agreement and divergence. We provide detailed insights in Section A2.
4 Experimental setup
We employ seven models to generate stories from prompts sourced from three distinct datasets, using six decoding strategies with varying hyperparameter configurations. To evaluate the quality of the generated text, we rely on three widely used automatic metrics: coherence, diversity, and MAUVE. In addition, human evaluators are engaged to provide further assessment of the quality of the generations.
4.1 Models
We employ GPT2-XL (1.5B parameters) (Radford et al., 2019), Mistral 7B v0.1 and v0.3 (Touvron et al., 2023), Llama 3.1 8B (Dubey et al., 2024), Deepseek 7B (DeepSeek-AI et al., 2024), Qwen2 7B (Yang et al., 2024a), and Falcon 2 11B (Malartic et al., 2024).
Models Datasets Metrics Decoding strategy Hyperparameter Values # Experiments
Deepseek Wikitext Coherence Beam search Beam width {3, 5, 10, 15, 20, 50} 7 x 3 x 6 = 126
Falcon2 Wikinews Diversity Contrastive search $k$1, 3, 5, 10, 15, 20, 50} 7 x 3 x 7 x 5 = 735
GPT2-XL Book MAUVE $\alpha$0.2, 0.4, 0.6, 0.8, 1.0}
Llama3 QText Adaptive contrastive search NA NA 1 x 3 x 1 = 3
Mistralv01 Human Evaluation Sampling with temperature Temperature {0.1, 0.3, 0.5, 0.7, 0.9, 1.0} 7 x 3 x 6 = 126
Mistralv03 Top $k$ sampling $k$1, 3, 5, 10, 15, 20, 50} 7 x 3 x 7 = 147
Qwen2 Top $p$ (nucleus) sampling $p$0.6, 0.7, 0.8, 0.9, 0.95} 7 x 3 x 5 = 105
Grand Total 1,242
Table 1: Overview of the experimental setup. The total number of experiments was determined by the combinations of models, datasets, and hyperparameter values explored. Falcon2 (11B) was run on an NVIDIA A100 GPU, while all other models were evaluated using an NVIDIA GeForce RTX 4090.
4.2 Hyperparameters
For contrastive search, we evaluate combinations of $\alpha\in\{0.2,0.4,0.6,0.8,1.0\}$ and $k\in\{1,3,5,10,15,20,50\}$, while for beam search, we consider beam width $\in\{3,5,10,15,20,50\}$. For sampling with temperature we consider temperature $\in\{0.1,0.3,0.5,0.7,0.9,1.0\}$, for top-$k$ sampling, we use $k\in\{1,3,5,10,15,20,50\}$ and for top-$p$ (nucleus) sampling we evaluate $p\in\{0.6,0.7,0.8,0.9,0.95\}$, for a total of 60 hyperparameter combinations. We also use the hyperparameter-free adaptive contrastive search for comparison, which increase the number of experiments by 3, totaling to 1242 experiments and 2.2 Mio generated stories. Further details are presented in Table 1.
4.3 Evaluation Metrics
We follow Su and Xu (2022) and use three metrics to automatically measure the quality the generations: Diversity, MAUVE, and Coherence.
Diversity.
This metric aggregates $\mathrm{n}$-gram repetition rates:
$\text{DIV}=\prod_{n=2}^{4}\frac{\mid\text{ unique }\mathrm{n}\text{-grams }\left(\mathrm{x}_{\text{cont }}\right)\mid}{\text{ total }\mathrm{n}\text{-grams }\left(\mathrm{x}_{\text{cont }}\right)\mid}$
A low diversity score suggests the model suffers from repetition, and a high diversity score means the model-generated text is lexically diverse.
MAUVE.
MAUVE (Pillutla et al., 2021) score measures the distribution similarity between the set of generated text and the set of gold references.
Coherence.
Proposed by Su et al. (2022), the coherence metric is defined as the averaged log-likelihood of the generated text conditioned on the prefix text as
$\operatorname{Coherence}(\hat{\boldsymbol{x}},\boldsymbol{x})=\frac{1}{|\hat{\boldsymbol{x}}|}\sum_{i=1}^{|\hat{\boldsymbol{x}}|}\log p_{\mathcal{M}}\left(\hat{\boldsymbol{x}}_{i}\mid\left[\boldsymbol{x}:\hat{\boldsymbol{x}}_{<i}\right]\right)$
where $\boldsymbol{x}$ and $\hat{\boldsymbol{x}}$ are the prefix text and the generated text, respectively; [:] is the concatenation operation and $\mathcal{M}$ is the OPT model (2.7B) Zhang et al. (2022). Finally, we apply a smoothed Min-Max normalization to the coherence values to ensure consistency and comparability with Diversity and MAUVE metrics.
$\operatorname{COH}=\frac{\text{Coherence}-\min(\text{Coherence})+1}{\max(\text{Coherence})-\min(\text{Coherence})+1}$
Aggregation.
QText
Following a generalization of the F1-score for three metrics, we use the harmonic mean of Diversity, MAUVE, and Coherence as aggregation measure:
$\text{QText}=\frac{3}{\frac{1}{\text{DIV}}+\frac{1}{\text{MAUVE}}+\frac{1}{\text{COH}}}$
Values close to one indicate high-quality text generation, while values approaching zero reflect low-quality outcomes.
In this study, we base our metric evaluations on multiple datasets, yielding a total of 1,242 measurements.
Human Eval.
In order to evaluate the quality of the generated text, we consider two critical aspects: fluency, coherence and similarity to human text. A fluent piece of text is written in grammatical English and has a natural flow (e.g. excluding unnatural repetition or web formatting). A coherent piece of text should stay on topic with the prompt and avoid unnatural topic drift. We provide five native English speakers with 39 competing continuations for three prompts and ask them to rate them based on their perceived quality. Definitions and instructions for the rating process are shown in Appendix A4, Figure 27.
4.4 Datasets
We evaluate our proposed method across three domains in open-ended text generation: news, Wikipedia, and stories. For the news domain, we utilize 2,000 articles from Wikinews; for the Wikipedia domain, we draw from the WikiText-103 dataset (1,314 examples; Merity et al., 2016); and for the stories domain, we employ the Project Gutenberg split of BookCorpus (1,947 examples; Zhu et al., 2015). Each example consists of a prompt paired with a gold-standard, human-generated continuation for evaluation purposes.
For each prompt, we generate 256 tokens as a continuation. The resulting text is assessed using both automatic metrics (outlined in Section 4.3 ) and human evaluations.
5 Results
Automatic Evaluation
Type Strategy Configuration Coherence Diversity MAUVE QText
Deterministic Beam search (Beam width) 3 -0.72 12.78 25.27 21.50
5 -0.69 12.14 22.01 19.92
10 -0.64 10.52 18.33 17.24
15 -0.62 9.40 16.37 15.59
20 -0.60 8.52 14.85 14.26
50 -0.55$\star$ 6.17 9.94 10.19
Contrastive search ($\alpha$, $k$) (0.2, 1) -0.75 9.90 24.21 18.25
(0.2, 3) -0.95 20.73 42.30 33.95
(0.2, 5) -0.98 22.75 44.22 36.44
(0.2, 10) -1.03 24.70 47.12 39.13
(0.2, 15) -1.07 25.82 47.85 40.32
(0.2, 20) -1.12 26.48 48.68 41.12
(0.2, 50) -1.28 29.18 52.06 44.38
(0.4, 1) -0.75 9.90 24.15 18.25
(0.4, 3) -1.24 42.11 59.86 54.02
(0.4, 5) -1.39 49.02 61.35 57.98
(0.4, 10) -1.61 55.31 62.70 59.77
(0.4, 15) -1.81 58.85 63.90 61.28
(0.4, 20) -1.96 61.62 63.44 62.15
(0.4, 50) -2.16 70.77 69.42 70.08
(0.6, 1) -0.78 10.68 25.54 19.60
(0.6, 3) -1.44 62.11 70.46 68.52
(0.6, 5) -1.62 69.89 73.76 72.76
(0.6, 10) -1.99 76.58 72.12 73.28
(0.6, 15) -2.29 78.88 66.85 70.21
(0.6, 20) -2.51 80.03 59.57 65.73
(0.6, 50) -3.39 84.08 29.30 38.70
(0.8, 1) -0.75 9.87 23.94 18.18
(0.8, 3) -1.97 77.12 74.94 74.07
(0.8, 5) -2.56 81.38 62.02 66.39
(0.8, 10) -3.51 84.52 29.78 40.97
(0.8, 15) -4.03 85.84 17.31 23.74
(0.8, 20) -4.34 86.77 14.52 18.48
(0.8, 50) -5.13 87.85 12.21 12.96
(1.0, 1) -0.75 9.85 24.05 18.18
(1.0, 3) -3.11 83.95 58.92 63.33
(1.0, 5) -3.93 85.42 29.09 37.63
(1.0, 10) -4.73 86.96 13.36 17.28
(1.0, 15) -5.03 87.51 11.51 12.76
(1.0, 20) -5.20 87.62 11.31 11.69
(1.0, 50) -5.76 87.96 11.28 10.85
Adaptive contrastive search - -1.68 93.94$\star$ 79.61 85.72$\star$
(Hyperparameter-free)
Stochastic Sampling with temperature ($t$) 0.1 -0.76 10.56 25.19 19.26
0.3 -0.92 17.16 37.28 28.97
0.5 -1.20 34.88 60.67 50.63
0.7 -1.70 66.07 82.85 75.78
0.9 -2.48 88.38 90.85 84.58
1 -3.07 93.27 91.17$\star$ 83.30
Top-$k$ ($k$) 1 -0.76 10.10 24.93 18.66
3 -1.41 50.04 72.75 63.86
5 -1.66 65.69 81.48 74.82
10 -1.90 77.10 85.80 80.28
15 -1.98 77.67 84.40 78.65
20 -2.05 79.20 84.21 78.86
50 -2.23 79.85 85.33 78.32
Top-$p$ (nucleus) ($p$) 0.6 -1.36 46.66 72.96 62.86
0.7 -1.56 61.38 80.80 73.60
0.8 -1.78 74.50 85.84 80.76
0.9 -2.05 84.09 89.18 84.50
0.95 -2.21 87.60 90.05 85.31
Reference Human - -2.74 93.28 100.00 87.37
Table 2: Automatic evaluation results for all seven models across three datasets. Scores are aggregated as weighted averages, accounting for the different number of generations per dataset. The highest scores for each strategy are highlighted in bold, with overall best results additionally marked by a $\star$.
In this section, we present our main results. Table 2 shows combinations of models, decoding strategies, and hyperparameter selections, along with their automatic evaluations in terms of Diversity, MAUVE, and Coherence. The values are computed as weighted means across the three datasets.
The results confirm the sensitivity of text generation quality to the specific values chosen for decoding strategies. The highest-ranking strategies consistently achieve a balance between coherence and diversity, demonstrating that successful text generation often hinges on not favoring one quality at the expense of the other. In contrast, lower-ranking strategies tend to overemphasize one aspect—either coherence or diversity—leading to less optimal outputs. For example, beam search typically favors coherence at the cost of diversity, while contrastive search with high values of $\alpha$ and $k$ prioritizes diversity, resulting in less coherent text.
Interestingly, the results also reveal that smaller models, such as GPT2-XL (1.5B), can outperform larger models like Falcon2 (11B) when the correct decoding strategy is applied. This finding underscores the importance of selecting optimal hyperparameters and decoding methods, which can significantly impact performance, regardless of model size.
In terms of human evaluation, there is a moderate correlation with automatic evaluation metrics, but the variability suggests that existing automatic metrics may not fully capture the nuanced preferences of human raters. As previously noted by Su and Collier (2023) and Garces-Arias et al. (2024), metrics like MAUVE tend to favor diverse text outputs, even when coherence decreases. Our findings confirm this observation, as human judgments often align more closely with a balance between coherence and diversity, rather than with the extremes favored by individual metrics. Nevertheless, as shown in Figures 2 and 3, there are similarities in the extremes of the spectrum: Both automatic scoring through QText and human scoring show preferences for decoding strategies that balance coherence and diversity better, such as contrastive search with moderate values of $\alpha$ and $k$, adaptive contrastive search, and sampling methods that allow diversity. A comprehensive overview of hyperparameter areas of high performance is presented in Table 2 and more detailed in Figures 4, 6, 19, 20, 22 and 24.
Finally, the moderate correlation between human evaluations and automatic metrics, as presented in Figure 26, underscores the need for more reliable and robust automatic evaluation methods. While current metrics offer valuable insights, they fall short of accurately reflecting the quality of open-ended text generation as perceived by human evaluators. This highlights the ongoing challenge of developing automatic metrics that can more effectively balance the competing factors of coherence and diversity in open-ended text generation tasks.
Figure 2: Top five and bottom five decoding strategies across datasets, based on human scores. The highest-ranking strategies generally strike a balance between coherence and diversity, while the lowest-ranking strategies tend to overemphasize one at the expense of the other—such as beam search, which favors coherence, or contrastive search with $\alpha=1.0$ and $k=50$, which prioritizes diversity.
Figure 3: Top five and bottom five decoding strategies across datasets, based on QText. The score exhibits similarities to human preferences in Figure 2.
6 Practical Recommendations
Based on the results of our study, we provide the following practical recommendations for selecting decoding strategies in open-ended text generation tasks. To achieve a balance between coherence and diversity, we recommend using contrastive search with moderate values of $\alpha$ and $k$. Specifically, setting $\alpha$ to 0.6 and $k$ to either 5 or 10 has consistently led to high-quality text outputs in our experiments. These parameter ranges strike an effective balance between generating coherent and diverse text, as visualized in the appendix. Our analysis shows a sharp performance decline when moving toward extreme values, such as low $\alpha$ or low $k$ (which lead to overly coherent but dull text), or high $\alpha$ and high $k$ (which produce diverse but incoherent text).
In addition to manually tuned hyperparameters, we recommend considering hyperparameter-free solutions, such as adaptive contrastive search, which dynamically adjusts the parameters during generation. This approach mitigates the need for fine-tuning specific values and has shown promising results in generating well-balanced text without requiring extensive parameter exploration.
On the other hand, beam search is generally not recommended for open-ended text generation. While beam search often excels in tasks that require high coherence, it tends to produce repetitive and less diverse text, making it unsuitable for creative or varied content generation.
Sampling methods, including top-k and nucleus (p) sampling, can perform well but tend to exhibit higher variance in output quality. Among these methods, we recommend using top-k sampling with $k=50$ and nucleus sampling with $p=0.9$ or $p=0.95$. These parameter settings have consistently produced high-quality text that aligns well with both automatic and human evaluation metrics. While these strategies are effective, their inherent randomness can sometimes lead to variability, so multiple generations may be necessary to achieve optimal results.
It is also important to caution against relying solely on automatic evaluation metrics when assessing text quality. Our results demonstrate that current automatic metrics are insufficient to fully capture human preferences. As noted in previous studies, including Garces-Arias et al. (2024), certain decoding methods, such as contrastive search with the DoubleExp method, achieve high scores on metrics like MAUVE but are consistently rejected by human evaluators for lack of coherence or clarity. This suggests that more research is needed to develop robust evaluation metrics that better correlate with human judgments of text quality.
Lastly, we encourage users to consider the specific dataset when selecting decoding strategies. For example, tasks like story generation (as in the book prompt) tend to reward creativity, so strategies that promote diversity may be preferable. In contrast, tasks like Wikipedia or news generation (as in Wikinews) demand more factual accuracy and coherence, where stricter control over diversity may be more beneficial. Detailed performance analyses for these different scenarios are provided in the appendix, offering further insights into which strategies work best for each dataset type.
7 Conclusion
Decoding strategies for large language models (LLMs) are a crucial yet often underexamined aspect of open-ended text generation. This study presents a detailed analysis of how hyperparameter selection across various decoding strategies significantly influences the quality of generated text. The findings highlight the necessity of maintaining a balance between coherence and diversity, as strategies that heavily prioritize one tend to underperform overall. Through extensive sensitivity analysis, we show that the choice of decoding method and its associated hyperparameters can impact text quality as much as, if not more than, model size.
We provide practical guidelines for decoding strategies, recommending balanced approaches such as contrastive search with moderate values of $\alpha$ and $k$, alongside hyperparameter-free methods like adaptive contrastive search. While sampling methods, such as top-k and nucleus sampling, can produce high-quality text, they exhibit greater variability and demand careful tuning.
Human evaluations demonstrate a moderate correlation with automatic metrics, pointing to the need for more reliable and robust evaluation tools. Current metrics sometimes prioritize diversity over coherence, which may not always align with human judgment. Our findings support the growing consensus that more nuanced metrics are required to better capture the trade-offs considered by human evaluators when assessing text quality. These insights aim to assist practitioners in understanding the critical role of hyperparameter selection in open-ended text generation.
Limitations
While this study provides insights into how hyperparameters affect the quality of generated text, it also has limitations. First, the study is confined to open-ended text generation tasks, and further research is necessary to assess the applicability of our methods to other NLP tasks such as summarization and machine translation. Additionally, the impact of our findings on models that have undergone supervised fine-tuning (SFT) remains unexplored. Future work could investigate how these models perform under similar experimental conditions.
Moreover, the analysis was conducted exclusively in English, leaving the generalizability to multilingual settings untested. Previous studies Su and Collier (2023); Garces-Arias et al. (2024) suggest that these findings may generalize to other languages, but further investigation is required to confirm this.
A key observation from our experiments is the discrepancy between individual automatic evaluation metrics and human judgments, highlighting the need for more reliable automatic evaluation methods. Given the resource-intensive nature of human evaluations, developing comprehensive automatic metrics that more accurately capture both coherence and diversity, while better aligning with human preferences, remains a critical area for future work.
Due to resource constraints, conducting an extensive human evaluation covering more combinations of models and decoding strategies was not feasible. This limitation underscores the necessity for automatic metrics that better correlate with human judgment, as a comprehensive human evaluation would be prohibitively costly.
Finally, this study focuses on analyzing the sensitivity of commonly used automatic metrics, which are inherently influenced by the models employed in the evaluation process. For example, coherence is assessed using the OPT 2.7B model, and MAUVE is calculated with GPT-2-based embeddings. The potential biases introduced by these models have not been fully explored, and addressing these biases in future research may offer valuable insights.
Ethics Statement
We affirm that our research adheres to the ACL Ethics Policy. This work involves the use of publicly available datasets and does not include any personally identifiable information. For our human evaluation, we employed third-party evaluators, ensuring a rate of over \$20 per hour. An ethical concern worth mentioning is the use of language models for text generation, which may produce harmful content, either through intentional misuse by users or unintentionally due to the training data or algorithms. We declare that there are no conflicts of interest that could potentially influence the outcomes, interpretations, or conclusions of this research. All funding sources supporting this study are acknowledged in the acknowledgments section. We have diligently documented our methodology, experiments, and results, and commit to sharing our code, data, and other relevant resources to enhance reproducibility and further advancements in the field.
Acknowledgments
Matthias Aßenmacher received funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) as part of BERD@NFDI, under grant number 460037581.
Appendix
Appendix A1 Case Studies
Prompt "The atmosphere dripped with emotion. Even Sola, the 13-year old, caught the current. Her eyes, fixated on the mother-son scene",
Beam search (10) "in front of her, welled up with tears.
’I’m so proud of you,’ the mother said.
’I’m so proud of you,’ the son said.
’I’m so proud of you,’ the mother said.
’I’m so proud of you,’ the son said… (text continues)"
Contrastive search (0.6, 10) "unfavor-
ablc: 0.5pt;
[…]
Contrastive search (1.0, 50) ", were moist.
’I’m sorry, Mama,’ the boy said.
’I’m sorry too, son,’ the mother replied.
’I’m sorry for not being a good son. I’m sorry for not being there for you… (text continues)"
Adaptive contrastive search ", shimmered with tears. A few minutes later, I was on the roof of the school, peering through binoculars to see what was happening down below. In the distance, I could see the silhouettes of a number of school buses, their windows blackened by fire… (text continues)"
Sampling with temperature (1) "unfolding before her, darted from one to the other, but she could not see it. Not really.
’You’re not my son,’ the mother spoke, her voice strong in spite of the tears rolling down her cheeks… (text continues)"
Top-k (50) "reflected a desire to be a mother. Her heart was full of love for her mother. She felt the mother’s pain. She had lost her own mother to cancer.
A 13-year-old, a 24-year-old, and a 40-year-old stood in the center of the church… (text continues)"
Top-p (0.9) "were shining with silent approval. She had seen something unusual. She had witnessed an act of parental care. The son was trying to calm his mother. He spoke words of assurance; he told her things would work out… (text continues)"
Human ", were wide with wonder.
’Mommy, why is he crying?’ she asked softly.
’It’s okay, Sola. It’s okay. Don’t worry,’ her mother replied… (text continues)"
Table 3: Case Study: Comparison of multiple decoding strategies for a prompt taken from the Book corpus. Beam search exhibits a repetitive behavior, while contrastive search generates erratic continuations for some of the hyperparameter choices.
Prompt "Two French journalists kidnapped on Wednesday north-east of the capital Kabul are believed to be in good health, although the condition of their three Afghan assistants is unknown.",
Beam search (10) "Two French journalists kidnapped on Wednesday north-east of the capital Kabul are believed to be in good health, although the condition of their three Afghan assistants is unknown.
Two French journalists kidnapped on Wednesday north-east of the capital Kabul are believed to be in good health, although the condition of their three Afghan assistants is unknown… (text continues)"
Contrastive search (0.6, 10) "Sylvain Dufoix and Nicolas Hénin of Canal+ were abducted by gunmen in broad daylight, along a road that connects Taloqan, the capital of Takhar province, 200 kilometers (125 miles). A spokesman for President Hamid Karzai condemned the ’brutality and barbarianism of terrorists who target civilians.’ … (text continues)"
Contrastive search (1.0, 50) "Reporters without Borders (RSF) condemns the ’barbaric’ abduction of the two reporters and their Afghan assistants, and calls on the Taliban to release them immediately.
’Two years after the murder of seven French journalists in Mali and five years after the killing of the American journalist James Foley by IS, the world’s media are once again being targeted with this barbaric kidnapping.’ … (text continues)"
Adaptive contrastive search "French President Nicolas Sarkozy said on Thursday that he had been in contact with the Taliban, who have claimed responsibility for the kidnapping.
’I have asked the Taliban to release the hostages,’ Sarkozy said. ’I have asked the Taliban to release the hostages unharmed.’ … (text continues)"
Sampling with temperature (1) "French President Nicolas Sarkozy has confirmed that two French journalists kidnapped in northern Afghanistan were in good health, although he rejected the Taliban’s call for the French government to withdraw its troops from Afghanistan.
’They are in good health,’ Sarkozy said, speaking in New Delhi. ’We are doing everything to ensure their return.’ … (text continues)"
Top-k (50) "The two French journalists, Nicolas Hénin and Edouard Elias, were kidnapped by gunmen in the Tagab district of Kapisa province, north-east of Kabul, on Wednesday.
Their Afghan assistants, who were also kidnapped, are believed to be in poor health, according to a statement from the French foreign ministry. … (text continues)"
Top-p (0.9) "The two French journalists, Nicolas Hénin and Edouard Elias, were kidnapped by gunmen in the Tagab district of Kapisa province, north-east of Kabul, on Wednesday.
Their Afghan assistants, who were also kidnapped, are believed to be in poor health, according to a statement from the French foreign ministry. … (text continues)"
Human "France’s foreign ministry said in a statement on Thursday that the three had been ’kidnapped by armed individuals who took them from their hotel in Herat province at gunpoint.’ Afghan President Ashraf Ghani’s spokesman Aimal Faizi said a helicopter had been sent to pick up the journalists … (text continues)"
Table 4: Case Study: Comparison of multiple decoding strategies for a prompt taken from the Wikinews corpus. Beam search exhibits a repetitive behavior.
Prompt "Headlam served as Officer Commanding North @-@ Western Area in 1946, and as Director of Training from 1947 to 1950. In 1950 2̆013 51,",
Beam search (10) "he was Commandant of the Royal Military Academy Sandhurst. From 1951 to 1953, he was Deputy Chief of the Imperial General Staff. From 1953 to 1955, he was Chief of the Imperial General Staff… (text continues)"
Contrastive search (0.6, 10) "Headlamexercised command responsi"
Contrastive search (1.0, 50) "he was the Officer Commanding the 1st Battalion, The Royal Ulster Rifles. He was promoted to Brigadier in 1951, and served as the Director of Army Education from 1951 to 1954… (text continues)"
Adaptive contrastive search "he was Commandant of the Royal Military Academy, Sandhurst, and from 1951 to 1953, he was Deputy Chief of the Imperial General Staff. From 1953 to 1955, he was Chief of the Imperial General Staff… (text continues)"
Sampling with temperature (1) "he served as Commander, 1st Battalion, Royal Canadian Regiment. In 1951–53, he was Chief of Staff of Canadian Army Staff, and in 1953–55, he was responsible for the reorganization of the Canadian Army Staff as Director of Operations… (text continues)"
Top-k (50) "he served as the Commanding Officer of the 1st Battalion of the Nigerian Regiment. Headlam was a member of the first intake of the Nigerian Military School in 1933. He attended the Royal Military College, Sandhurst in 1935… (text continues)"
Top-p (0.9) "he was Commandant of Depot @-@ Training, and the next year was Deputy Chief of @-@ Staff Training. Headlam returned to the Far East, serving as Commander, Australian Troops on the Malay Peninsula from 1952 to 1954… (text continues)"
Human "he was Director of Operations and Intelligence, and in 1951–54, Commander of the 1st Division, which was the most powerful division in the world. He was appointed Commander-in-Chief of the Army in 1954… (text continues)"
Table 5: Case Study: Comparison of multiple decoding strategies for a prompt taken from the Wikitext corpus. Beam search exhibits a repetitive behavior, while contrastive search and sampling-based strategies generate erratic continuations for some of the hyperparameter choices.
Appendix A2 Performance per Model, Dataset and Decoding Strategy
Figure 4: Distribution of metric values per model, by using a Beam Search decoding strategy.
Figure 5: Effect of beam width on metric behavior. Falcon2 is used for visualization, as the trends observed were consistent across all models.
Figure 6: Distribution of metric values per model, by using a Contrastive Search decoding strategy.
Figure 7: Effect of $\alpha$ (fix) and $k$ on metric behavior (model Deepseek visualized).
Figure 8: Effect of $k$ (fix) and $\alpha$ on metric behavior (model Deepseek visualized).
Figure 9: Effect of $\alpha$ (fix) and $k$ on metric behavior (model Falcon2 visualized).
Figure 10: Effect of $k$ (fix) and $\alpha$ on metric behavior (model Falcon2 visualized).
Figure 11: Effect of $\alpha$ (fix) and $k$ on metric behavior (model GPT2-XL visualized).
Figure 12: Effect of $k$ (fix) and $\alpha$ on metric behavior (model GPT2-XL visualized).
Figure 13: Effect of $\alpha$ (fix) and $k$ on metric behavior (model Llama3 visualized).
Figure 14: Effect of $k$ (fix) and $\alpha$ on metric behavior (model Llama3 visualized).
Figure 15: Effect of $\alpha$ (fix) and $k$ on metric behavior (model Mistralv03 visualized).
Figure 16: Effect of $k$ (fix) and $\alpha$ on metric behavior (model Mistralv03 visualized).
Figure 17: Effect of $\alpha$ (fix) and $k$ on metric behavior (model Qwen2 visualized).
Figure 18: Effect of $k$ (fix) and $\alpha$ on metric behavior (model Qwen2 visualized).
Figure 19: Distribution of metric values per model, by using an Adaptive Contrastive Search decoding strategy, here we report results for GPT2-XL only.
Figure 20: Distribution of metric values per model, by using a Sampling with temperature decoding strategy.
Figure 21: Effect of temperature on metric behavior, visualized by model.
Figure 22: Distribution of metric values per model, by using a Top-$k$ Sampling decoding strategy.
Figure 23: Effect of k on metric behavior, visualized by model.
Figure 24: Distribution of metric values per model, by using a Top-$p$ Nucleus Sampling decoding strategy.
Figure 25: Effect of p on metric behavior, visualized by model.
Appendix A3 Correlation of Human evaluation and Automatic metrics
Figure 26: Correlation between human scores and automatic metrics. Moderate correlations are observed between human judgments and automatic metrics such as QText and MAUVE, while Coherence and Diversity show weak to very weak correlations.
Appendix A4 Human Evaluation Form
Figure 27: Human Evaluation form, including general instructions and definitions for the evaluation criteria.